# HMT: Hierarchical Memory Transformer for Long Context Language Processing

### v2

@@ Abstract
Transformer-based large language models (LLM) have been widely used in language processing applications. However, most of them restrict the context window that permits the model to attend to every token in the inputs. Previous works in recurrent models can memorize past tokens to enable unlimited context and maintain effectiveness. However, they have “flat” memory architectures, which have limitations in selecting and filtering information. Since humans are good at learning and self-adjustment, we speculate that imitating brain memory hierarchy is beneficial for model memorization. We propose the Hierarchical Memory Transformer (HMT), a novel framework that enables and improves models’ long-context processing ability by imitating human memorization behavior. Leveraging memory-augmented segment-level recurrence, we organize the memory hierarchy by preserving tokens from early input token segments, passing memory embeddings along the sequence, and recalling relevant information from history. Evaluating general language modeling (Wikitext-103, PG-19) and question-answering tasks (PubMedQA), we show that HMT steadily improves the long-context processing ability of context-constrained and long-context models. With an additional 0.5% $\sim$ 2% of parameters, HMT can easily plug in and augment future LLMs to handle long context effectively. Our code is open-sourced on Github: https://github.com/OswaldHe/HMT-pytorch.
Segment-level Recurrence, Large Language Model, Memory Hierarchy
@@ 1 Introduction
Transformer (Vaswani et al., 2017) has demonstrated its strength in contextual learning and is utilized in various applications in language processing (Dong et al., 2019) and computer vision (Dosovitskiy et al., 2020). For a decoder-only transformer model, each transformer block contains a self-attention and a feedforward network module. An optimized self-attention layer has a quadratic computational and linear space complexity (Dao et al., 2022) regarding the sequence length since it computes interactions between each token and all previous tokens in the sequence. To maintain the inference speed and satisfy memory requirements, most transformer models enforce maximum sequence length. For example, the Llama 2 model is designed to process 4096 tokens at maximum (Touvron et al., 2023). However, real-world applications involving long documents, such as book summarization (Rae et al., 2019) and lifelong question-answering tasks (Sun et al., 2019; Dai et al., 2022), can have an extensive number or even infinite stream of inputs.
@@
Existing research attempts to build long-range transformers using sparse attention (Beltagy et al., 2020; Zhang et al., 2021; Kitaev et al., 2020), retrieval-augmented models (Bertsch et al., 2023; Wu et al., 2022), and recurrent sequence models (Peng et al., 2023; Gu & Dao, 2023; Rae et al., 2019). However, these models have at least one of two issues: (1) difficult to train or adapt to other models due to a change in the core model architecture and (2) low effectiveness for long-range inputs with frequent context switching. In this work, we propose the Hierarchical Memory Transformer (HMT), a novel framework to enable and enhance the long-context language processing ability of any model. HMT transforms any pretrained model into a memory-augmented segment-level recurrent model that imitates the brain’s memory hierarchy and human memorization behavior. It has the unique features as follows:
@@
Hierarchical Memorization: HMT mimics the memory hierarchy of the brain (Burgin, 2011) employing both learned memory tokens and past input tokens. HMT stratifies memory into sensory, short-term, and long-term, with interactions between each other.
@@
Memory Recall Mechanism: HMT imitates memory recall by storing encoded memory embeddings generated from previous iterations and searching based on the relevance to current token segments. It also handles inputs switching between multiple relevant topics.
@@
One key advantage of using HMT over other memory-augment models is that HMT is a model-independent plug-and-play framework: future decoder-only models can directly use HMT without extra implementations. With joint training and fine-tuning of newly introduced parameters and original parameters of the backbone model, evaluation exhibits HMT’s advantage when applied to various models. Our impacts include:
  • HMT applies to any models with a short and limited context window to enable learning infinitely long context with enhanced effectiveness  [^1] In this paper, effectiveness means how well the language model generates meaningful texts. It is measured in perplexity., especially with context switching. We demonstrate that with HMT, OPT and OpenLlamaV2 can be 25.5% and 17.6% more effective in perplexity (PPL) on Wikitext-103 with multiple samples concatenated, and 11.4% and 9.48% on PG-19, a book dataset. We also evaluated HMT with Llama 2 7B regarding the PubMedQA dataset with multiple contexts. It was 9.81% more effective in long-answer contextual reasoning with a 1.0% higher short-answer prediction accuracy.
  • With a similar size state-of-the-art backbone model, HMT outperforms various models designed for long context processing. Moreover, HMT can further increase the effectiveness of these long-context models. For recurrent sequence models such as RWKV, HMT with OpenLlamaV2 is 44.0% and 14.9% more effective on Wikitext-103 and PG-19 respectively. Applying HMT to RWKV can further enhance effectiveness by 16.5% on Wikitext-103 and 9.96% on PG-19.
  • HMT is more effective than Recurrent Memory Transformer (RMT), which is the previous state-of-the-art framework for augmenting models with a long context processing ability. HMT outperforms RMT by 13% for the Wikitext-103 dataset and 5.42% for the PG-19 dataset in PPL. HMT is also 8.98% more effective than RMT for long-answer contextual reasoning and 1.8% more accurate in short-answer prediction over QA tasks, indicating the advantage in long-context question-answering tasks.
@@ 2 Related Works and Problem Formulation
2.1 Long-range Transformer
@@
Since one of the bottlenecks of transformers is the quadratic computational complexity of self-attention, a natural approach is sparsifying attention computation. A naive sparse attention pattern is the sliding window attention (Kovaleva et al., 2019), where each token attends to neighbors within a local window. However, this neglects long-range interaction between words. Existing works such as Longformer (Beltagy et al., 2020) and Poolingformer (Zhang et al., 2021) extend the sliding window attention by adding global attending tokens and applying pooling to increase the receptive field area. Unlimiformer (Bertsch et al., 2023) adopts the retrieval-augmented generative model by searching the top K most important tokens for the incoming sequence and applying attention in the decoders to just those tokens, which prunes computations with minor loss. Nevertheless, the contribution of less relevant tokens may accumulate over time and impact the overall sequence generation. Although these methods extend the attainable context length, they cannot prevent increasing memory consumption as the input length increases. Alternatively, compressing past tokens using a recurrent sequence model can potentially reduce memory consumption by compressing information into a fixed-size embedding.
Figure 1: Overall workflow of HMT. (1) For a given segment of tokens, HMT will first perform representation extraction utilizing the segment summarization prompt embedding ($H^{T}$) to summarize part of the segment. (2) The generated segment summary embedding ($H_{n}^{sum}$) is used with the cached memory embeddings for memory search with cross attention. The output is a memorization prompt embedding ($H^{S}_{n}$) which contains information relevant to the current segment and can guide the summarization in Step 4. (3) The memorization prompt embedding and the last k embeddings from the previous segment will augment the segment. (4) The backbone model (BBM) will process the augmented segment and generate hidden embeddings for logits ($H_{n}^{out}$) and the memory embedding of the $n^{th}$ segment ($H_{n}^{mem}$), which will be pushed into the cache containing previous memory embeddings.
2.2 Recurrent Sequence Model
@@
Recurrent Neural Networks (RNN) have been extensively explored in sequence processing research, including Long Short-term Memory (Hochreiter & Schmidhuber, 1997) and Gated Recurrent Unit (Chung et al., 2014). They reveal strong performance in memorizing past information and are hardware-friendly for implementing customized accelerators (Chang et al., 2015). However, RNNs have limited advantages compared with self-attention in learning contextual relationships between words in language processing (Bahdanau et al., 2014). One approach to alleviate this issue is the coarse-grain recurrence, in which the model splits inputs into segments, performs attention inside each segment, and propagates states between segments. The Compressive Transformer (Rae et al., 2019) further stores and compresses previous states to enhance memorization. The Recurrent Memory Transformer (RMT) (Bulatov et al., 2022) utilizes a memory token to summarize and propagate segment information without modifying the transformer block architecture. Theoretically, they can process unlimited long sequences, but previous information will be diluted after multiple summarizations and generation quality can drop when less relevant information occupies the memory. Nevertheless, it provides an important building block for HMT.
@@
$\text{BBM}(*)$ Backbone model
$H^{S}_{n}$ Memorization prompt embedding
@@
$H_{n}^{mem}$ memory embedding of the $n^{th}$ segment
$D$ Discarded embeddings
@@
$H^{mem}_{[n-N+1,n)}$ Cached memory embeddings
$H^{T}$ Segment summarization prompt embedding
$H_{n}^{sum}$ Summary embedding of the $n^{th}$ segment
Another approach augments RNN that involves interactions between the current inputs and the previous states to learn contextual relationships in a similar way as self-attention and accelerate the computation with linear convolution. One of the representatives, RWKV (Peng et al., 2023), is an RNN model inspired by the attention-free transformer (AFT) (Zhai et al., 2021). It includes a time-mixing module to learn from previous states and a channel-mixing module to learn from the previous output. Mamba (Gu & Dao, 2023) is a state-space model that employs gated convolution to accelerate the inference of recurrence. These models are energy and memory-efficient with fast training speed but have limitations on recalling information from a very large document.
2.3 Problem Formulation: Adaptive Long-context Processing
We aim to develop a model that processes long-context information adaptively with the following properties:
Continuous learning: The model can process an infinite stream of data and gradually improve the performance when processing similar data.
Context Adaptability: Based on the context/topic of the input stream, the model can adaptively select past relevant information to enhance effectiveness, since irrelevant context can distract the model (Shi et al., 2023).
In real-world applications, restrained by memory bandwidth and capacity, as well as data generation speed, long documents cannot be read as a whole by the computing hardware (Agerri et al., 2015). Furthermore, users who are constantly interacting with the language model can refer to the previous topic or switch to another topic that has high relevance to past information provided. To be effective, most recurrent models need to encode all previous inputs in the states, which can contain irrelevant information that will degrade the model’s effectiveness.
3 HMT Method
Table 1 describes all notations we use to illustrate the HMT architecture in this section.
3.1 Overall Workflow
HMT processes input in units of segments of $L$ tokens. For every segment, HMT walks through 4 steps as shown in Figure 1:
@@ 3.1 Overall Workflow
    Representation extraction, which encodes partial segment as a single embedding.
@@
    Memory search, which utilizes the generated embedding as a query to find relevant information in the history.
@@
    Prepending sensory memory, which augments the segment to capture tokens in the previous segment and relevant history.
@@
    Segment processing and summarization, which processes the augmented segment to get hidden embeddings for generating logits and a memory embedding that summarizes the augmented segment.
@@
The first two steps are the memory recall mechanism discussed in Section 3.2. Steps 3 and 4 are explained in Section 3.3 along with the concept of hierarchical memorization.
3.2 Memory Recall Mechanism
@@
To handle context switching and prevent the intervention of irrelevant context, HMT performs memory recall to extract related information from past knowledge. The memory recall mechanism involves three steps: representation extraction, memory search, and memory augmentation.
@@
Representation Extraction: Depicted in Step 1 of Figure 1, HMT selects the first $j$ embeddings from the hidden embeddings of the $n^{th}$ segment, $H_{n}$, to extract the topic of the segment. The embeddings are augmented with the segment summarization prompt embedding $H^{T}$. $H^{T}$ is a learnable parameter and is identical for every segment, deployed to prompt the backbone model (BBM) to summarize the segment from the first $H^{T}$ to the second occurrence of $H^{T}$. Instead of extracting from the token embedding of BBM, we use a learnable embedding to allow a larger prompt embedding space for summarization. The backbone model will then process the augmented embeddings and generate a new embedding at the end of the output as the representation of the segment:
$D\circ H^{sum}_{n}=\text{BBM}([H^{T}\circ H_{n}[0,j)\circ H^{T}])$
@@
where $D$ is the embedding that will be discarded, $H^{sum}_{n}$ is the summary embedding of the $n^{th}$ segment only, and $\text{BBM}(*)$ represents the backbone model. It will be used for memory search. $\circ$” is the concatenation operator.
@@
Memory Search: Presented in Step 2 of Figure 1, $H^{sum}_{n}$ is utilized as a query to find relevant memory embeddings generated from Step 4 when processing previous segments. We store the last $N$ embeddings ($H^{mem}_{[n-N+1,n)}$) and then computes:
$Q=H^{sum}_{n}W_{q},K=H^{mem}_{[n-N+1,n)}W_{k}$
$H^{S}_{n}=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{h}}})H^{mem}_{[n-N+1,n)}$
@@
where $d_{h}$ is the hidden dimension of the cross attention. The computation is similar to cross-attention without value and output projection. $\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{h}}})$ calculates the normalized similarity score, and applying it directly to $H^{mem}_{[n-N+1,n)}$ can ensure similar distributions of output value and old memory tokens. We expect that the projection $W_{q}$ and $W_{k}$ can be trained such that summarizations containing similar contexts have high attention scores after projections. In practice, it performs better than having the value and output projections.
@@
The output of a memory search is a memorization prompt embedding $H^{S}_{n}$ containing information relevant to the $n^{th}$ segment. It will be applied to augment the $n^{th}$ segment.
Notice that HMT’s memory is accumulative: the $n^{th}$ memory embedding contains information of all previous $n-1$ segments, with a higher loss of information for older segments. Recalling memory aims at strengthening the relevant memory by rolling back to relevant topics.
The overhead of the memory recall mechanism can be hidden if implemented on configurable hardware, as explained in Appendix G.

### v3

# HMT: Hierarchical Memory Transformer for Efficient Long Context Language Processing

@@ Abstract
Transformer-based large language models (LLM) have been widely used in language processing applications. However, due to the memory constraints of the devices, most of them restrict the context window. Even though recurrent models in previous works can memorize past tokens to enable unlimited context and maintain effectiveness, they have “flat” memory architectures. Such architectures have limitations in selecting and filtering information. Since humans are good at learning and self-adjustment, we believe that imitating brain memory hierarchy is beneficial for model memorization. Thus, we propose the Hierarchical Memory Transformer (HMT)  [^1] https://github.com/OswaldHe/HMT-pytorch, a novel framework that facilitates a model’s long-context processing ability by imitating human memorization behavior. Leveraging memory-augmented segment-level recurrence, we organize the memory hierarchy by preserving tokens from early input segments, passing memory embeddings along the sequence, and recalling relevant information from history. Evaluating general language modeling, question-answering tasks, and the summarization task, we show that HMT consistently improves the long-context processing ability of existing models. Furthermore, HMT achieves a comparable or superior generation quality to long-context LLMs with 257×2$\sim$ 57$\times$2 ∼ 57 × fewer parameters and 2.5116×2.5$\sim$ 116$\times$2.5 ∼ 116 × less inference memory, significantly outperforming previous memory-augmented models.
HMT: Hierarchical Memory Transformer for Efficient Long Context Language Processing
Zifan He 1, Yingqi Cao 2, Zongyue Qin 1, Neha Prakriya 1, Yizhou Sun 1, and Jason Cong 1 1 University of California, Los Angeles, 2 University of California, San Diego zifanhe1202@g.ucla.edu, yic033@ucsd.edu, {qinzongyue, nehaprakriya, yzsun, cong}@cs.ucla.edu
@@ 1 Introduction
Transformer Vaswani et al. (2017) has demonstrated its strength in contextual learning and is utilized in various applications in language processing Dong et al. (2019) and computer vision Dosovitskiy et al. (2020). For a decoder-only transformer model, each transformer block contains a self-attention and a feedforward network module. An optimized self-attention layer has a quadratic computational and linear space complexity Dao et al. (2022) regarding the sequence length since it computes interactions between each token and all previous tokens in the sequence. To maintain the inference speed and satisfy memory requirements, most transformer models enforce maximum sequence length. For example, the Llama 3 model is designed to process 8192 tokens Dubey et al. (2024) and the Llama 2 can process up to 4096 tokens Touvron et al. (2023). However, real-world applications involving long documents, such as book summarization Rae et al. (2019) and lifelong question-answering tasks Sun et al. (2019); Dai et al. (2022), can have an enormous or even infinite stream of inputs.
@@
Existing research attempts to build long context transformers using sparse attention Beltagy et al. (2020); Zhang et al. (2021); Kitaev et al. (2020), retrieval-augmented models Bertsch et al. (2023); Wu et al. (2022), and recurrent sequence models Peng et al. (2023a); Gu and Dao (2023); Rae et al. (2019). Still, these models face at least one of two issues: (1) difficulty in adapting to future models due to a change in the core model architecture and (2) low effectiveness for long-range inputs under frequent context switching. In this work, we propose the Hierarchical Memory Transformer (HMT), a novel framework to enable and augment models’ long-context processing ability. HMT transforms models into a memory-augmented recurrent model that imitates the brain’s memory hierarchy and human memorization behavior. It has the following unique features:
@@
Hierarchical Memorization: HMT mimics the memory hierarchy of the brain Burgin (2011) employing both learned memory tokens and current input tokens. HMT stratifies memory into sensory, short-term, and long-term, with interactions between each other.
@@
Memory Retrieval Mechanism: HMT imitates memory recall by storing encoded memory embeddings generated from previous iterations and searching based on the relevance to current token segments.
@@
One key advantage of utilizing HMT over other memory-augment models is that HMT is a model-independent plug-and-play framework: future decoder-only models can directly serve as the backbone model of HMT to augment their long context processing ability without extra implementation efforts. With joint training and fine-tuning of newly introduced and original parameters of the backbone model, HMT is applicable to a wide range of LLMs, including transformer-based models and state-space models. Our contributions include:
  • HMT consistently improves models’ generation quality with long context for various model architectures. We demonstrate HMT on both transformer-based architecture and state-space models. Evaluating on Wikitext-103, PG-19 Rae et al. (2019), and PubMedQA Jin et al. (2019) datasets with multiple contexts concatenated, HMT can improve the effectiveness by up to 25.5% in perplexity and 1.0% higher prediction accuracy over the baseline models.
  • HMT with small backbone models can outperform large models trained on longer context samples, implying a high memory efficiency. We evaluate HMT with SmolLM Allal et al. (2024), OPT Zhang et al. (2022), and OpenLlamaV2 Geng and Liu (2023) models on the LongBench Bai et al. (2023b) benchmark. In sum, HMT can achieve comparable or higher metric results with 257×2$\sim$ 57$\times$2 ∼ 57 × fewer parameters and 2.5116×2.5$\sim$ 116$\times$2.5 ∼ 116 × lower inference memory requirement than long-context large language models.
  • HMT surpasses previous methods specialized for efficient long-context processing by compressing contexts. We compare HMT with RMT Bulatov et al. (2022), LongMem Wang et al. (2024), Memorizing Transformer Wu et al. (2022), CCM Kim et al. (2023), and HOMER Song et al. (2024), which are recent SoTA of memory-augmented and hierarchical methods. With the same or similar size backbone model, HMT has a better generation quality in both general language modeling and QA tasks. Furthermore, HMT has a lower memory complexity, indicating better scalability as the input length increases.
@@ 2 Related Works and Problem Formulation
2.1 Long Context Transformers
@@
Since one of the bottlenecks of transformers is the quadratic computational complexity of self-attention, a natural approach is sparsifying attention computation. A naive sparse attention pattern is the sliding window attention Kovaleva et al. (2019), where each token attends to neighbors within a local window. However, this neglects long-range interaction between words. Existing works such as Longformer Beltagy et al. (2020) and Poolingformer Zhang et al. (2021) extend the sliding window attention by adding global attending tokens and applying pooling to expand the receptive field area. Unlimiformer Bertsch et al. (2023) adopts the retrieval-augmented generative method by searching the top K most important tokens for the incoming sequence. It then applies attention to just those tokens in the decoders, resulting in pruned computations with minor losses. Nevertheless, the contribution of less relevant tokens may accumulate over time and impact the overall sequence generation. Although these methods extend the attainable context length, they cannot prevent increasing memory consumption as the input length increases. Alternatively, compressing past tokens using a recurrent sequence model can potentially reduce memory consumption by condensing the information into a fixed-size embedding.
2.2 Recurrent Sequence Models
@@
Recurrent Neural Networks (RNN) have been extensively explored in sequence processing research, including Long Short-term Memory Hochreiter and Schmidhuber (1997) and Gated Recurrent Unit Chung et al. (2014). They reveal that RNNs perform well in memorizing past information and are hardware-friendly for implementing customized accelerators Chang et al. (2015). However, RNNs have limited advantages in learning contextual relationships between words compared with self-attention in language processing Bahdanau et al. (2014). One approach to alleviate this issue is the coarse-grain recurrence, in which the model splits inputs into segments, performs attention inside each segment, and propagates states (i.e., compressed information as embeddings) between segments. The Compressive Transformer Rae et al. (2019) further stores and compresses previous states to enhance memorization. The Recurrent Memory Transformer (RMT) Bulatov et al. (2022) utilizes a memory token to summarize and propagate segment information without modifying the transformer block architecture. Theoretically, they can process unlimited long sequences, but previous information will be diluted after multiple summarizations and generation quality can drop when less relevant information occupies the memory. Recent works Chevalier et al. (2023); Kim et al. (2023) aim to further optimize RMT to improve generation quality by concatenating the results of summarizations, but this sacrifices the inference memory efficiency.
Another approach augments RNN by involving interactions between the current inputs and the previous states to learn contextual relationships in a similar way as self-attention and accelerate the computation with linear convolution. One of the representatives, RWKV Peng et al. (2023a), is an RNN model inspired by the attention-free transformer (AFT) Zhai et al. (2021). It includes a time-mixing module to learn from previous states and a channel-mixing module to learn from the previous output. Mamba Gu and Dao (2023) is another recurrent method based on the state-space model that employs gated convolution to accelerate model inference. These models are energy and memory-efficient with fast training speed and are able to achieve high performance in memorization tasks (e.g., associative recall), but have limitations on capturing contextual relationships and filtering irrelevant information. Recent works combine transformers with Mamba Lieber et al. (2024); Team et al. (2024) to mitigate this issue, but this reintroduces the scaling issue of the transformers.
2.3 Problem Formulation: Adaptive Long-context Processing
Figure 1: Overall workflow of HMT. For a segment, (1) HMT will first perform representation encoding, utilizing the segment summarization prompt embedding ($T$) to summarize part of the segment. (2) The generated segment summary embedding ($S_{n}$) is used with the cached memory embeddings for memory search with cross attention. The output is a memorization prompt embedding ($P_{n}$) which contains information relevant to the current segment. (3) The memorization prompt embedding and the last $k$ embeddings from the previous segment will augment the segment. (4) The backbone model (BBM) will process the augmented segment and generate hidden embeddings for logits ($H_{n}^{out}$) and the memory embedding ($M_{n}$), which will be pushed into the long-term memory.
We intend to develop a model that can handle infinitely long context inputs with context adaptability: Based on the context/topic of the input stream, the model can adaptively select past relevant information to enhance effectiveness, since irrelevant context can distract the model Shi et al. (2023).
In real-world applications, restrained by memory bandwidth and capacity, as well as data generation speed, long documents cannot be read as a whole by the computing hardware Agerri et al. (2015). Furthermore, users who are constantly interacting with the language model can refer to the previous topic or switch to another topic that has high relevance to past information. For effectiveness, most recurrent models need to encode all previous inputs in the states, which can contain irrelevant information and degrade the model’s quality.
3 Hierarchical Memory Transformer
The main idea of HMT is to store information hierarchically and search for relevant information throughout the memory hierarchy. Table 1 describes all notations we use to illustrate the HMT architecture in this section.
3.1 Overall Workflow
Given a backbone model to enhance, HMT chunks the input into $L$-token segments and operates on the hidden embeddings of the token segments ($\{H_{n}\}_{n=0}^{\infty}$), generated by the token embedding layer of the backbone model. For every segment $n$, HMT walks through four steps shown in Figure 1:
@@ 3.1 Overall Workflow
$\text{BBM}(\cdot)$ Backbone model
$P_{n}$ Memorization prompt embedding
@@
$M_{n}$ memory embedding of the $n^{th}$ segment
@@
$M_{[n-N+1,n)}$ Cached memory embeddings
$T$ Segment summarization prompt embedding
$S_{n}$ Summary embedding of the $n^{th}$ segment
@@
    Representation encoding by the backbone model, which encodes part of the segment containing the essence of the ongoing topic into a single embedding to represent its context, denoted by $H_{n}$.
@@
    Memory search, which utilizes the current context as a query to find relevant information in the memory.
@@
    Prepending sensory memory, which augments the segment to capture information in the previous segment and other relevant information.
@@
    Decoding and summarization, which processes the augmented segment to get hidden embeddings for generating logits and a memory embedding that summarizes the augmented segment.
@@
The first two steps are the memory retrieval mechanism discussed in Section 3.2. Steps 3 and 4 are explained in Section 3.3 along with the concept of hierarchical memorization.
3.2 Memory Retrieval Mechanism
@@
To handle context switching and prevent the intervention of irrelevant context, HMT performs memory retrieval to extract only relevant information from past knowledge. The memory retrieval mechanism involves three steps: representation extraction, memory search, and memory augmentation.
@@
Representation Encoding: Depicted in Step 1 of Figure 1, HMT selects the first $j$ embeddings from the hidden embeddings of the $n^{th}$ segment, $H_{n}$, to extract the topic of the segment. The embeddings are augmented with the segment summarization prompt embedding $\mathbf{T}$. $T$ is a learnable parameter embedding, deployed to prompt the backbone model (BBM) to summarize the segment by soft prompt tuning Liu et al. (2023). Instead of extracting from the token embedding of BBM, we make $T$ learnable to allow a larger prompt embedding space for summarization. The backbone model will then process the augmented embeddings and generate a new embedding at the end of the output as the representation of the segment:
$S_{n}=\text{BBM}([T||H_{n}[0,j)||T])[j,j+1)$ (1)
@@
where $S_{n}$ is the summary embedding of the $n^{th}$ segment only, $\text{BBM}(\cdot)$ is the backbone model, and “||||| |" is the concatenation operator. $S_{n}$ will be used for memory search.
@@
Memory Search: Shown in Step 2 of Figure 1, $S_{n}$ is utilized as a query to find relevant memory embeddings generated from Step 4 when processing previous segments. We keep a sliding window of $N$ embeddings ($M_{[n-N+1,n)}$) and then compute:
$Q_{n}=S_{n}W_{q},K_{n}=M_{[n-N+1,n)}W_{k}$ (2)
$P_{n}=\text{softmax}(\frac{Q_{n}K_{n}^{T}}{\sqrt{d_{h}}})M_{[n-N+1,n)}$ (3)
@@
where $d_{h}$ is the hidden dimension of the cross attention. The computation is similar to cross-attention without value and output projection. $\text{Softmax}(\frac{Q_{n}K_{n}^{T}}{\sqrt{d_{h}}})$ calculates the normalized similarity score and applies it directly to $M_{[n-N+1,n)}$ to ensure similar distributions of output value and old memory tokens. We expect that the projection $W_{q}$ and $W_{k}$ can be trained such that summarizations containing similar contexts have high attention scores after projections.
@@
The output of a memory search is a memorization prompt embedding $P_{n}$ containing information relevant to the $n^{th}$ segment. It will be applied to augment the $n^{th}$ segment. Notice that HMT’s memory is accumulative: the $n^{th}$ memory embedding contains information of all previous $n-1$ segments, with a higher loss of information for older segments. We hope that retrieving memory will strengthen the relevant memory and reduce this loss.
In practice, representation encoding is executed in parallel with the model inference on GPUs since they are independent tasks. Memory search has time complexity $O(N)$, and can also run in parallel with the segment inference when $N$ is small (e.g., $N=300$). Thus, the overall runtime overhead of HMT is negligible.